OpenStudy (anonymous):

help please! solve for y: 3^y+1= 9^y-1 1. 1 2. 2 3. 3 4. 4

OpenStudy (anonymous):

hi :DDD

OpenStudy (tkhunny):

You have written \(3^{y} + 1 = 9^{y} - 1\) You will have grave difficulty solving this. Please remember your Order of Operations and use parentheses to write what you mean. \(3^{y+1} = 9^{y-1}\) \(3^{y+1} = \left(3^{2}\right)^{y-1}\) \(3^{y+1} = 3^{2y-2}\) \(y+1 = 2y-2\)